# Studying magnetic circular vortex dichroism effect for photoionization of Rydberg atoms with vortex photons

### v1

@@ III.4 MCVD with different photon TAM $m_{\gamma}$
The asymtotic MCVD convergence to $\pm 1$ for $|m_{\gamma}|\geq 2$ is a common feature. This behavior reveals a fundamental chiral asymmetry in the photoionization dynamics of Rydberg atoms with vortex photons: At fixed TAM modulus, for Rydberg states with positive magnetic quantum number (${m_{l}}>0$), the ionization cross section is significantly enhanced when interacting with vortex photons carrying positive TAM quantum number ($m_{\gamma}\geq+2$) compared to those with negative ones ($m_{\gamma}\leq-2$). The parity-conserving interaction rules governing this process show a complementary preference for negative-$m$ states to couple more strongly with negative-$m_{\gamma}$ photons.
III.5 Discussions
We present a qualitative analysis of the observed phenomena, with circular Rydberg atoms (CRAs) serving as our primary example. While similar analysis can be extended to general Rydberg atoms, the latter case involves greater complexity.
III.5.1 Multiple expansion
We can perform a multiple expansion for the photon field, whereby Eq. (1) becomes
$\mathcal{M}_{fi}^{P}$ $=-N_{B}\frac{e}{m_{e}}\sum_{\sigma=0}^{\infty}\Bigg{[}\frac{1}{\sigma!}\int d^{3}\mathbf{r}\,e^{-i\mathbf{p}\cdot\mathbf{r}}(i\mathbf{k}\cdot\mathbf{r})^{\sigma}$
$=-N_{B}\frac{e}{m_{e}}\sum_{\sigma=0}^{\infty}\left[\frac{1}{\sigma!}(\mathbf{\epsilon}_{k{\Lambda_{\gamma}}}\cdot\mathbf{p})(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{\sigma}\tilde{\psi}_{i}(\mathbf{p})\right],$ (21)
where
[…]
The transition matrix for the vortex case is
$\mathcal{M}^{V}_{fi}$ $=N_{B}\int\frac{d\phi_{\mathbf{k}}}{2\pi}e^{im_{\gamma}\phi_{\mathbf{k}}}\mathcal{M}^{P}_{fi}(\mathbf{k})$
$=-N_{B}\frac{e}{m_{e}}\sum_{\sigma=0}^{\infty}\Bigg{[}\frac{1}{\sigma!}\int\frac{d\phi}{2\pi}e^{im_{\gamma}\phi_{\mathbf{k}}}(\mathbf{\epsilon}_{k{\Lambda_{\gamma}}}\cdot\mathbf{p})$
$\quad\times(-\mathbf{k}\cdot\mathbf{\nabla}_{p})^{\sigma}\tilde{\psi}_{i}(\mathbf{p})\Bigg{]}$
$=-N_{B}\frac{e}{m_{e}}e^{im_{\gamma}\phi_{\mathbf{p}}}\sum_{\sigma=0}^{\infty}\Bigg{[}\frac{1}{\sigma!}\int\frac{d\phi}{2\pi}e^{im_{\gamma}\phi}(\mathbf{\epsilon}_{k{\Lambda_{\gamma}}}\cdot\mathbf{p})$
$\quad\times(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{\sigma}\tilde{\psi}_{i}(\mathbf{p})\Bigg{]}.$ (23)
By considering Eq. (II.1), (III.5.1), and (III.5.1), we can analyze the dependence of multiple contributions on TAM of the vortex photon. For the $\sigma$-th order multiple contribution:
  • The factor $(-\mathbf{k}\cdot\mathbf{\nabla}_{p})^{\sigma}$ contains $\phi$-dependent phase terms with $m_{\nabla}=0,~\pm 1,~\cdots,~\pm\sigma$
The global $\phi$-dependent phase factor $e^{i(m_{\gamma}+m_{\Lambda}+m_{\nabla})\phi}$ serves as a filter when integrated over one complete cycle. This integration selects only terms satisfying the conservation condition: $m_{\gamma}+m_{\Lambda}+m_{\nabla}=0$, which constrains the photon TAM to $m_{\gamma}=-m_{\Lambda}-m_{\nabla}$, with allowed values spanning 0 to $\pm(\sigma+1)$. The multipole expansion exhibits the following characteristic hierarchy:
  • Dipole($\sigma=0$): $m_{\gamma}\in\{0,\pm 1\}$
  • Quadrupole($\sigma=1$): $m_{\gamma}\in\{0,\pm 1,\pm 2\}$
This pattern extends to higher multiple orders. For photo-ionization with vortex photon at fixed $m_{\gamma}$, the minimal contributing order is $\sigma=|m_{\gamma}|-1$ for $|m_{\gamma}|\geq 1$, while the $m_{\gamma}=0$ case begins with $\sigma=0$.
III.5.2 Approximations for CRAs
Since the operator $\partial_{p}$ always yields a $1/p$ term when acting on $\tilde{\psi}_{i}$, we can extract a global factor $k/p$ from Eq. III.5.1. This factor is typically small for non-relativistic photoionization processes. Consequently, higher-order terms generally has smaller contributions to Eq. III.5.1.
With the $\phi$ integral, we explicitly get $m_{\nabla}=-m_{\gamma}-m_{\Lambda}$. The lowest order contribution has $\sigma=|m_{\nabla}|=|m_{\gamma}+m_{\Lambda}|$. Eq. III.5.1 contains $\phi$-independent terms that has no contribution to the lowest order term, due to the filtering effect of $\phi$ integral. We can therefore neglect these terms when evaluating Eq. III.5.1, leading to the simplified expression:
[…]
For Rydberg atoms, the wave function in momentum space is given by Eq. 25:
$\tilde{\psi}_{nl{m_{l}}}(\mathbf{p})\propto\frac{Y_{l}^{m_{l}}(\theta_{\mathbf{p}},\phi_{\mathbf{p}})}{n^{2}a^{2}p^{l+4}}.$ (26)
[…]
$(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{\sigma}$ $\approx\left(\frac{k\sin\theta_{\mathbf{k}}}{p}\right)^{\sigma}\prod_{j=1}^{\sigma}\Bigg{[}\left(l+\frac{j+3}{2}\right)e^{-i\lambda\phi}$
$\quad+\left(\frac{j+3}{2}\right)e^{i\lambda\phi}\Bigg{]}.$ (27)
The higher-order contributions are strongly suppressed by the $(k\sin\theta_{\mathbf{k}}/p)^{\sigma}$ factor, with the $\sin\theta_{\mathbf{k}}$ term providing additional suppression.
III.5.3 phenomenon analyzing
As established earlier, the $\phi$ integral in Eq. III.5.1 acts as a selection filter through the condition $m_{\gamma}+m_{\Lambda}+m_{\nabla}=0$, where $m_{\Lambda}=0$ or $\pm 1$ originates from the polarization term:
The phase contributions $m_{\nabla}$ emerge from the operator $(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{\sigma}$.
For $|m_{\gamma}|\geq 2$, the dipole contribution ($\sigma=0$) vanishes, and the dominant contribution arises from $\sigma=|m_{\nabla}|\geq 1$. According to Eq. III.5.2, four distinct cases emerge:
$(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{|m_{\nabla}|}$ $\rightarrow\left(\frac{k\sin\theta_{\mathbf{k}}}{p}\right)^{|m_{\nabla}|}\prod_{j=0}^{|m_{\nabla}|-1}\left(l+2+\frac{j}{2}\right)$
$\quad\times e^{-i|m_{\nabla}|\phi}$ (29)
$(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{|m_{\nabla}|}$ $\rightarrow\left(\frac{k\sin\theta_{\mathbf{k}}}{p}\right)^{|m_{\nabla}|}\prod_{j=0}^{|m_{\nabla}|-1}\left(2+\frac{j}{2}\right)$
$\quad\times e^{-i|m_{\nabla}|\phi}$ (30)
$(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{|m_{\nabla}|}$ $\rightarrow\left(\frac{k\sin\theta_{\mathbf{k}}}{p}\right)^{|m_{\nabla}|}\prod_{j=0}^{|m_{\nabla}|-1}\left(l+2+\frac{j}{2}\right)$
$\quad\times e^{i|m_{\nabla}|\phi}$ (31)
$(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{|m_{\nabla}|}$ $\rightarrow\left(\frac{k\sin\theta_{\mathbf{k}}}{p}\right)^{|m_{\nabla}|}\prod_{j=0}^{|m_{\nabla}|-1}\left(2+\frac{j}{2}\right)$
$\quad\times e^{i|m_{\nabla}|\phi}$ (32)
The significant difference in prefactors between these cases explains why CRAs with large $l$ exhibit strong directional preference: photoionization favors vortex photons co-rotating ($m_{\gamma}>0$ for ${m_{l}}>0$, $m_{\gamma}<0$ for ${m_{l}}<0$) with the atomic angular momentum. This manifests in MCVD values approaching $\pm 1$ (Figs. 2, Fig. 5). This reveals the chiral nature in photoionization of polarized Rydberg atoms using vortex photons.
For $|m_{\gamma}|=1$, the dipole contribution ($\sigma=0$) plays important role. Two scenarios occur:
1. When $m_{\gamma}$ and $\Lambda_{\gamma}$} share the same sign, the $p_{\perp}\cos^{2}(\theta_{\mathbf{k}}/2)e^{-i{\Lambda_{\gamma}}\phi}$ term dominates, i.e. dipole contribution dominates, yielding MCVD $\approx 0$ resembling initial plane wave case (see $m_{\gamma}=1$ case in Figs. 1 and 2).
2. When $m_{\gamma}$ and $\Lambda_{\gamma}$} have opposite signs, the $p_{\perp}\sin^{2}(\theta_{\mathbf{k}}/2)e^{i{\Lambda_{\gamma}}\phi}$ term, which gives dipole contribution, is suppressed by $\sin^{2}(\theta_{\mathbf{k}}/2)$. $\sigma=1,2$ contributions (given by $p_{z}\sin\theta_{k}$ and $p_{\perp}\cos^{2}(\theta_{\mathbf{k}}/2)e^{-i{\Lambda_{\gamma}}\phi}$ terms respectively) become important. Their interference with dipole contribution introduce significant $m_{l}$}-dependent MCVD (see $m_{\gamma}=-1$ case in Figs. 1 and 2).
For $m_{\gamma}=0$, the dipole contribution is given by $p_{z}\sin\theta_{k}$ term and suppressed by $\sin\theta_{k}$. The quadrupole contribution ($\sigma=1$) given by $p_{\perp}\cos^{2}(\theta_{\mathbf{k}}/2)e^{-i{\Lambda_{\gamma}}\phi}$ term can be comparable to the dipole contribution. Their interference introduce significant $m_{l}$}-dependent MCVD (see $m_{\gamma}=0$ case in Figs. 1 and 2).
@@ IV Conclusions
Figures 1 and 2 demonstrate that the magnetic circular vortex dichroism (MCVD) is highly sensitive to the magnetic moment of Rydberg atoms when the incident vortex photons carry specific low values of total angular momentum (TAM). This sensitivity establishes optical vortices as a powerful tool for probing quantum magnetism in polarized Rydberg systems, paralleling their established role in studying optical magnetism in chiral materials [28]. Thus, vortex-based spectroscopy offers a novel approach for characterizing the magnetic properties of Rydberg atoms.
Experimentally, the MCVD of Rydberg atoms can be measured by irradiating the sample with identical vortex photon beams along the atomic polarization axis from opposite directions. The reversal of the vortex photons’ longitudinal momentum does not affect the measurement outcomes. Moreover, the energy of the initial vortex photon must be carefully selected to observe significant physical phenomena, as demonstrated in Figures 3 and 4.
We also show that he photoionization process exhibits strong angular momentum selectivity, preferentially favoring aligned configurations between photon TAM and atomic magnetic moment. The enhanced cross sections for co-rotating systems versus counter-rotating configurations provide definitive evidence of this angular momentum matching. This pronounced asymmetry directly reflects the chiral nature of vortex photon-Rydberg atom interactions.
@@ Appendix A Calculation of ${\cal M}^{V}$ with residue theorem
$\quad\times\sum_{\sigma=0}^{l}\sum_{{\mu=\mu_{\rm min}}}^{{\mu_{\rm max}}}f_{l{m_{l}},\sigma\mu}(\mathbf{p})(X_{a}^{+}+X_{b}^{+}+X_{c}^{+}),$ (34)

### v2

@@ III.4 MCVD with different photon TAM $m_{\gamma}$
The asymptotic convergence of MCVD to $\pm 1$ is a characteristic feature for $|m_{\gamma}|\geq 2$. As demonstrated in Figs. 4 and 5 for small $|m_{\gamma}|$ values (e.g., 2 and 3), and confirmed in Fig. 8 for larger values, this behavior is consistently maintained across all integer values of $|m_{\gamma}|\geq 2$. Theoretical analysis has been given after Figs. 4 and 5. This behavior reveals a fundamental chiral asymmetry in the photoionization dynamics of Rydberg atoms with vortex photons: At fixed TAM modulus, for Rydberg states with positive magnetic quantum number (${m_{l}}>0$), the ionization cross section is significantly enhanced when interacting with vortex photons carrying positive TAM quantum number ($m_{\gamma}\geq+2$) compared to those with negative ones ($m_{\gamma}\leq-2$). The parity-conserving interaction rules governing this process show a complementary preference for negative-$m$ states to couple more strongly with negative-$m_{\gamma}$ photons.
@@ IV Conclusions
This paper examines the photoionization of hydrogen-like Rydberg alkali atoms subjected to vortex photon beams. We propose a novel effect termed magnetic circular vortex dichroism (MCVD) and investigate its behavior across various scenarios. MCVD serves as a key physical observable that provides insights into the characteristics of hydrogen-like Rydberg alkali atoms.
As shown in Figures 4 and 5, MCVD exhibits high sensitivity to the magnetic moment of Rydberg atoms when the incident vortex photons carry specific low values of total angular momentum (TAM). This sensitivity establishes optical vortices as a powerful tool for probing quantum magnetism in polarized Rydberg systems, paralleling their established role in studying optical magnetism in chiral materials [1]. It is important to note that the energy of the incident vortex photon must be carefully chosen to observe significant physical phenomena, as demonstrated in Figures 6 and 7. Furthermore, we show that the photoionization process exhibits strong angular momentum selectivity, preferentially favoring aligned configurations between photon TAM and atomic magnetic moment. The enhanced cross sections for co-rotating systems versus counter-rotating configurations provide clear evidence of this angular momentum matching. This pronounced asymmetry directly reflects the chiral nature of vortex photon-Rydberg atom collisions.
Experimentally, the MCVD of Rydberg atoms can be measured by irradiating the sample with identical vortex photon beams along the atomic polarization axis from opposite directions. The reversal of the vortex photons’ longitudinal momentum does not affect the measurement outcomes.
@@ Appendix A Calculation of ${\cal M}^{V}$ with residue theorem
$\quad\times\sum_{{{\eta}}=0}^{l}\sum_{{\mu=\mu_{\rm min}}}^{{\mu_{\rm max}}}f_{l{m_{l}},{{\eta}}\mu}(\mathbf{p})(X_{a}^{+}+X_{b}^{+}+X_{c}^{+}),$ (38)